Grade 11General Physics

From a tower of height H, a particle is thrown vertically upwards with a speed u. The time taken by the particle to hit the ground is n time that taken by it to reach the highest to its path. Shoe that 2gH=nu²(n-2)

Jessy
8 Years agoGrade 11

1 Answer

Chirag Vaswani
8 Years ago
Time taken by the particle to reach highest point:
 
v = u – gt
at peak v = 0
u = gt
t1= u/g
 
to reach ground
s = ut – ½gt2
s = – H
- H = ut – ½gt2
½ gt2 – ut – H =0
 
t2 = u/g + (u2/g2 + 2H/g)½
discard the negative solution as time cannot be negative
t2 = u/g + (u2/g2 + 2H/g)½
t= t+ (t12 + 2H/g)½ 
t2 = nt1 
(nt- t1)2 =  t1+ 2H/g
(n – 1)u2/g2 = u2/g+2H/g
(n2 – 2n +1) u= u+ 2Hg
n(n-2)u2 = 2Hg
QED